Home   >  Worksheets   >  GCSE questions by topic   >  Probability equation questions

Probability equation questions

Free Printable Probability equation questions PDF Worksheets, Revision Guide & GCSE Past Paper Questions (Grade 8+)

Download free probability equation questions worksheets (PDF), GCSE maths past paper questions and worked solutions for Probability equation questions (8L). This is a Grade 8+ probability topic from the GCSE questions by topic collection for Foundation and Higher tier GCSE Maths (Edexcel, AQA, OCR and IGCSE).

The free PDF worksheets table below brings together printable question booklets (Qs) and worked answers or mark schemes (Ans) from Maths Genie, Dr Frost, Dr Austin, E Classroom and PMT, along with past paper question booklets.

Below the worksheet links you can read our probability equation questions revision notes and worked examples with diagrams, try real GCSE past paper probability equation questions questions with step-by-step solutions, or download a PDF of all probability equation questions worksheet links.

Free Probability equation questions PDF Worksheets & Answers

Grade 8+ • Topic 8L
Resource Questions Answers
GCSE past paper questions & solutions
Qs Ans
Graded exam questions & solutions
Qs Ans
GCSE worksheet & solutions
Qs Ans
Practice grid / worksheet & answers
Qs Ans
Practice grid / worksheet & answers
Qs Ans
GCSE topic booklet & answers
Qs Ans
PMT (Edxl)
Edexcel GCSE topic questions & solutions
Qs Ans
PMT (AQA)
AQA GCSE topic questions & solutions
Qs Ans
PMT (IGCSE)
Edexcel IGCSE topic questions & solutions
Qs Ans
GCSE Maths Grade 8+ Revision Notes & Method

Guide to Probability equation questions: Explanation, Diagrams & Worked Examples

Download Printable One Page Guide (PDF)

Explanation:

Algebraic probability questions give the number of counters in terms of n or x (or a ratio) and state the numerical probability of a two-step event without replacement. Write the branch probabilities as algebraic fractions, remembering to subtract 1 from the total denominator on the second pick.

Multiply (and add if there are multiple valid paths), equate your algebraic fraction to the given probability, cross-multiply to clear fractions, and rearrange into a quadratic equation an2 + bn + c = 0. Always reject negative or non-integer roots.

Example 1: A bag has n red counters and 3 blue counters. Two are taken without replacement. Given P(both red) = 514, find n.

n/(n+3) 3/(n+3) R B (n−1)/(n+2) 3/(n+2) R → 5/14 B

Solution:

nn + 3 × n − 1n + 2 = 514 ⇒ 14n(n − 1) = 5(n + 3)(n + 2)
3n2 − 13n − 10 = 0 ⇒ (3n + 2)(n − 5) = 0 ⇒ Answer: n = 5

Example 2: A bag has x green and (10 − x) yellow balls. Two are picked without replacement. Given P(both green) = 13, find x.

Solution:

Total balls = 10: P(GG) = x10 × x − 19 = 13
x(x − 1)90 = 13 ⇒ x2 − x = 30
x2 − x − 30 = 0 ⇒ (x − 6)(x + 5) = 0
Answer: x = 6
Scanned GCSE Exam Questions & Worked Solutions

GCSE Past Paper Probability equation questions Questions & Answers

Past Paper Example 1 IGCSE Jan 2020 2HR Q16 2 marks
Cody has two bags of counters, bag **A** and bag **B**.
Each of the counters has either an odd number or an even number written on it.
There are 10 counters in bag **A** and 7 of these counters have an **odd** number written on them.
There are 12 counters in bag **B** and 7 of these counters have an **odd** number written on them.
Cody is going to take at random a counter from bag **A** and a counter from bag **B**.
(a) Complete the probability tree diagram. [2 marks]
Bag ABag B7/10.............OddnumberEvennumber..........................Odd numberEven number..........................Odd numberEven number
Worked Solution & Mark Scheme (2 marks)
Part (a) Solution:
Bag A Even = 3/10 M1
Bag B branches = 7/12 Odd, 5/12 Even A1
Bag ABag B7/103/10OddnumberEvennumber7/125/12Odd numberEven number7/125/12Odd numberEven number
Answer (a): Bag A: 3/10; Bag B: 7/12, 5/12 and 7/12, 5/12
Past Paper Example 2 IGCSE Jun 2023 1FR Q28 5 marks
Moeen has a biased 6-sided dice.
The table gives information about the probability that, when the dice is thrown, it will land on each number.
NumberProbability1x20.1530.54y50.1360.03
(a) Show that x + y = 0.19 [2 marks]
(b) work out the value of x and the value of y
Show clear algebraic working.
[3 marks]
Worked Solution & Mark Scheme (5 marks)
Part (a) Solution:
1 – (0.15 + 0.5 + 0.13 + 0.03) = 1 – 0.81 = 0.19 M1 C1
Answer (a): x + 0.15 + 0.5 + y + 0.13 + 0.03 = 1 ⇒ x + y + 0.81 = 1 ⇒ x + y = 0.19
Part (b) Solution:
Adding equations: 4x = 0.28 ⇒ x = 0.07 M2
y = 0.19 – 0.07 = 0.12 A1
Answer (b): x = 0.07, y = 0.12
Past Paper Example 3 GCSE Jun 2018 3F Q24 5 marks
There are some counters in a bag.
The counters are red or white or blue or yellow.
Bob is going to take at random a counter from the bag.
The table shows each of the probabilities that the counter will be blue or will be yellow.
redwhiteblue0.45yellow0.25ColourProbability
There are 18 blue counters in the bag.
The probability that the counter Bob takes will be red is twice the probability that the counter will be white.
(a) Work out the number of red counters in the bag. [4 marks]
(b) Explain why. [1 mark]
Worked Solution & Mark Scheme (5 marks)
Part (a) Solution:
P(red) + P(white) = 1 – (0.45 + 0.25) = 0.30 ⇒ P(red) = 0.20, P(white) = 0.10 M2
Total counters = 18 ÷ 0.45 = 40 M1
Red counters = 0.20 × 40 = 8 A1
Answer (a): 8
Part (b) Solution:
Because half (0.5) of the total number of marbles must be a whole number
Answer (b): Because half (0.5) of the total number of marbles must be a whole number

← 8K Vector proof

Back to GCSE questions by topic

8M Invariant points →